OpenStudy (hansolo321):

9x – 4y = 43 2x + 12y = 79 solve by substitution with work and explanation pleasee. WILL FAVORITE AND AWARD MEDAL

OpenStudy (qwert2):

First you want to rework one of the two equations so that you can solve for x Next plug in the new equation into the equation that you did not rework. For example, say that x=3y. That would mean you plug 3y into where x would be in the equation. After you've done that, you should have y. Plug in that number into one of the equations to get x. Does that make sense?

OpenStudy (hansolo321):

@qwert2 this is one question from my homework. can you work it out for me so i can use it as example

OpenStudy (qwert2):

Ok I'll try. The first step would be to solve 2x+12y=79 for x.  once you got that, plug in -6y+39.5 into 9x-4y=43 like this: 9(-6y+39.5) -4y=43. Solve for y.